# TeleMem: Building Long-Term and Multimodal Memory for Agentic AI

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks [5, 32]. However, their effectiveness in long-term interactions remains fundamentally constrained by the context window limitations inherent to Transformer architectures. Even though contemporary long-context models can process hundreds of thousands of tokens [11], simply enlarging the context window does not fully address the issue: as interaction rounds accumulate, models struggle to allocate attention to distant tokens, making it difficult to accurately retrieve user-specific information embedded deep in the dialogue history. As a result, they may overlook key details or even forget important personalized information over time [12], significantly hindering their practical utility in scenarios requiring sustained, long-term relationships with users.
@@
This inherent limitation has motivated the development of more advanced memory architectures capable of structured storage, organization, and dynamic updating of information [31]. For example, MemoryBank [39] introduces a decay mechanism inspired by the Ebbinghaus forgetting curve to selectively preserve salient information. Mem0 [2] proposes a meaning-aware memory framework in which the LLM itself extracts atomic facts, retrieves semantically related memories, and decides whether to add, update, delete, or ignore them in a streaming, turn-by-turn process. Rsum [21] employs hierarchical summarization to distinguish short-term details from long-term abstractions, while Zep [18] represents dialogue history as a temporal knowledge graph to capture causal and temporal dependencies. Collectively, these systems extend RAG by incorporating notions of forgetting, abstraction, and structure, offering more principled approaches for managing long-term conversational memory.
@@ 3 Experiments
We conduct experiments on two ultra-long dialogue datasets to assess agent memory capabilities. ZH-4O [1] serves as a Chinese role-playing benchmark comprising 28 authentic human-LLM sessions (avg. 600 turns), annotated with 1,068 multiple-choice probing questions to test memory recall.
@@
Our evaluation strategies are tailored to each dataset: for ZH-4O [1], we quantify memory fidelity via QA Accuracy, where the model selects the single correct option for each multiple-choice query against ground-truth labels.
3.1 Results
ZH-4O
Table 1 reports the QA Accuracy of all methods on the ZH-4O benchmark. Our proposed TeleMem achieves the highest performance at 86.33%, surpassing the strong Long context LLM baseline (84.92%) that leverages the full dialogue history without explicit memory management. This demonstrates that TeleMem not only effectively compresses and retrieves salient information from ultra-long interactions but also mitigates the noise and redundancy inherent in raw conversation logs.
Among the specialized memory systems, Memobase (76.78%) performs best, likely due to its structured user profiling that aligns well with the role-playing nature of ZH-4O. In contrast, retrieval-augmented approaches such as RAG (62.45%) lag significantly, indicating that simple semantic retrieval without temporal or relational awareness is insufficient for multi-turn memory-intensive tasks. Notably, all memory-augmented methods outperform RAG, underscoring the necessity of explicit memory structures for long-horizon dialogue understanding. The consistent advantage of TeleMem over both agentic (A-Mem) and modular (Mem0) architectures further suggests that its read-write coordination mechanism enables more precise and contextually grounded memory access.
@@ 4 Related Work
Parametric memory aims to encode knowledge or contextual information directly into the weights or persistent hidden states of neural networks, distinguishing itself from non-parametric approaches that rely on external vector databases. Wang et al. [23] propose a decoupled memory mechanism, employing a residual side network to cache long-term context while keeping the base LLM frozen. Similarly, MLPMemory [26] internalizes retrieval by training multilayer perceptrons to approximate k-nearest neighbor distributions as differentiable mappings. Pushing the internalization of retrieval further, Tay et al. [19] introduce the Differentiable Search Index, which eliminates external indices by training the model to map queries directly to document identifiers via its parameters. regarding personalized adaptation, Zhang et al. [36] combine LoRA with Bayesian optimization to inject dialogue history into model weights. Complementary to these injection methods, Meng et al. [14] explore the interpretability of parametric storage, proposing ROME to locate and directly edit factual associations within the Transformer’s MLP layers. Overall, parametric memory methods offer scalability and unified reasoning but face challenges regarding update costs and potential misalignment with the base model.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks [5, 37, 29]. However, their effectiveness in long-term interactions remains fundamentally constrained by the context window limitations inherent to Transformer architectures. Even though contemporary long-context models can process hundreds of thousands of tokens [10], simply enlarging the context window does not fully address the issue: as interaction rounds accumulate, models struggle to allocate attention to distant tokens, making it difficult to accurately retrieve user-specific information embedded deep in the dialogue history. As a result, they may overlook key details or even forget important personalized information over time [12], significantly hindering their practical utility in scenarios requiring sustained, long-term relationships with users.
@@
This inherent limitation has motivated the development of more advanced memory architectures capable of structured storage, organization, and dynamic updating of information [40, 36]. For example, MemoryBank [45] introduces a decay mechanism inspired by the Ebbinghaus forgetting curve to selectively preserve salient information. Mem0 [2] proposes a meaning-aware memory framework in which the LLM itself extracts atomic facts, retrieves semantically related memories, and decides whether to add, update, delete, or ignore them in a streaming, turn-by-turn process. Rsum [25] employs hierarchical summarization to distinguish short-term details from long-term abstractions, while Zep [22] represents dialogue history as a temporal knowledge graph to capture causal and temporal dependencies. Collectively, these systems extend RAG by incorporating notions of forgetting, abstraction, and structure, offering more principled approaches for managing long-term conversational memory.
@@ 3 Experiments
We conduct experiments on two ultra-long dialogue datasets to assess agent memory capabilities. ZH-4O [1] serves as a Chinese role-playing benchmark comprising 28 authentic human-LLM sessions (avg. 600 turns), annotated with 1,068 multiple-choice probing questions to test memory recall. Complementarily, we utilize LoCoMo [16] as an English dataset, which consists of 10 long-context sessions paired with 1,540 question-answer pairs covering single-hop, multi-hop, open-domain, and temporal reasoning. Notably, we exclude the LoCoMo adversarial subset due to the absence of ground-truth answers.
@@
Our evaluation strategies are tailored to each dataset: for ZH-4O [1], we quantify memory fidelity via QA Accuracy, where the model selects the single correct option for each multiple-choice query against ground-truth labels. For LoCoMo [16], following prior work [2], we employ an LLM-as-a-Judge ($J$) approach using GPT-4o [5], which assesses the factual accuracy, relevance, and completeness of the generated responses.
3.1 Results on ZH-4O
@@
Main Results
Table 1 presents the performance of different memory paradigms on the ZH-4O benchmark, revealing clear distinctions in their ability to support long-horizon, multi-turn question answering. We make the following observations.
  • Retrieval-only methods are insufficient. RAG achieves the lowest accuracy (62.45%), indicating that flat semantic retrieval without temporal ordering or relational structure fails to support multi-turn, memory-intensive reasoning required by ZH-4O.
  • Explicit memory mechanisms consistently improve performance. All memory-augmented approaches outperform RAG, demonstrating the necessity of maintaining persistent and updatable memory states for long-horizon dialogue understanding. Among them, Memobase performs best (76.78%), likely due to its structured user profiling that aligns well with the role-playing characteristics of the benchmark.
  • Long-context modeling alone has inherent limitations. The Long Context LLM baseline achieves 84.92% accuracy by leveraging the full dialogue history, but its reliance on raw context makes it susceptible to noise, redundancy, and attention dilution as interactions grow longer.
  • Coordinated read–write memory yields the best results. Our proposed TeleMem attains the highest accuracy of 86.33%, outperforming both long-context and prior memory-based architectures. This demonstrates that selectively compressing salient information and enabling context-aware memory access is more effective than unstructured context accumulation or loosely coupled memory modules.
Write $\downarrow$     Read $\rightarrow$ 0.6B 1.7B 4B 8B 14B 32B Avg.
0.6B 52.72 64.23 73.69 75.47 74.06 72.19 68.73
1.7B 61.14 71.54 79.12 78.56 80.81 77.72 74.82
4B 63.48 74.25 81.09 83.33 84.55 83.33 78.34
8B 65.54 77.15 84.36 86.33 85.77 84.93 80.68
14B 67.04 78.28 84.55 86.33 87.55 85.49 81.54
32B 68.91 80.81 83.99 86.70 86.61 85.96 82.16
Avg. 63.14 74.38 81.13 82.79 83.17 81.60
Table 2: Memory read–write scaling law. Performance (QA Accuracy, %) when using different model sizes for the memory write LLM (rows) and the memory read LLM (columns). Results show consistent performance gains as either component scales up, with larger read–write combinations yielding the strongest performance.
Read-Write Scaling Law
Table 2 reveals a clear memory read–write scaling law by varying the model sizes of the memory write LLM and the memory read LLM. Overall performance improves monotonically as either component scales up, indicating that both memory writing and memory reading contribute substantially to downstream QA accuracy. However, the gains are not symmetric across the two dimensions. Increasing the size of the write LLM leads to consistent improvements across almost all read settings, suggesting that stronger writers produce more informative, compact, and robust memory representations that benefit readers of different capacities. Scaling the read LLM yields even more pronounced gains, particularly when paired with medium-to-large write models. This indicates that memory retrieval and reasoning capacity plays a critical role in exploiting stored information, especially when the memory content is sufficiently well-structured. Notably, performance saturates when the reader significantly outscales the writer, implying that retrieval capacity alone cannot compensate for low-quality memory writing. The best results are achieved when both the write and read LLMs are scaled jointly, highlighting the complementary and interdependent nature of memory writing and reading. These findings suggest that effective long-term memory systems should be designed with coordinated read–write capacity rather than over-optimizing either component in isolation.
3.2 Results on LoCoMo
Main Results
Table 3 compares different methods on the LoCoMo benchmark across four reasoning categories. Overall, the Long Context LLM achieves the strongest performance (70.71%), demonstrating the effectiveness of processing full interaction histories for complex memory-centric tasks. This advantage is particularly pronounced on temporal questions, where access to complete chronological context is crucial. In contrast, retrieval-based baselines perform poorly across all categories, highlighting the limitations of static semantic retrieval for long-term conversational reasoning. Memory systems exhibit diverse strengths across reasoning types. Mem0 performs well on single-hop questions (54.96%) and temporal reasoning (60.28%), suggesting that its modular memory abstraction is effective for fact recall and time-sensitive information. Memobase excels on multi-hop reasoning (66.04%), likely benefiting from its structured memory representations that better support compositional inference. A-Mem shows more balanced but moderate performance, indicating that agentic memory control alone is insufficient without strong memory structuring. TeleMem achieves competitive performance on single-hop (64.53%) and temporal reasoning (78.47%), approaching the Long Context LLM in these categories, which suggests that its memory abstraction effectively captures salient factual and temporal information. However, its weaker performance on multi-hop questions indicates remaining challenges in supporting complex relational reasoning over stored memories. Overall, these results highlight that different memory mechanisms favor different reasoning patterns, and no single approach uniformly dominates all categories, underscoring the importance of task-aware memory design for long-horizon QA.
Method Single-Hop Multi-Hop Open Domain Temporal Overall
RAG 20.91 32.39 35.41 48.03 39.03
Mem0 54.96 31.15 40.62 60.28 52.01
A-Mem 44.32 35.82 33.33 56.59 48.57
Memobase 50.70 66.04 37.50 58.97 57.59
Long context LLM 64.89 47.66 42.70 84.66 70.71
TeleMem 64.53 20.56 40.62 78.47 61.49
Table 3: Performance comparison on the LoCoMo benchmark. All metrics represent accuracy (%) evaluated via LLM-as-a-Judge across 1,540 question-answer pairs.
@@ 4 Related Work
Parametric memory aims to encode knowledge or contextual information directly into the weights or persistent hidden states of neural networks [14, 13, 18], distinguishing itself from non-parametric approaches that rely on external vector databases. Wang et al. [27] propose a decoupled memory mechanism, employing a residual side network to cache long-term context while keeping the base LLM frozen. Similarly, MLPMemory [31] internalizes retrieval by training multilayer perceptrons to approximate k-nearest neighbor distributions as differentiable mappings. Pushing the internalization of retrieval further, Tay et al. [23] introduce the Differentiable Search Index, which eliminates external indices by training the model to map queries directly to document identifiers via its parameters. regarding personalized adaptation, Zhang et al. [42] combine LoRA with Bayesian optimization to inject dialogue history into model weights. Complementary to these injection methods, Meng et al. [17] explore the interpretability of parametric storage, proposing ROME to locate and directly edit factual associations within the Transformer’s MLP layers. Overall, parametric memory methods offer scalability and unified reasoning but face challenges regarding update costs and potential misalignment with the base model.